OpenStudy (anonymous):

derivative of y = sin2x cos2x

OpenStudy (anonymous):

write it as y= sin4x/2 , then find the derivative using chain rule :)

OpenStudy (anonymous):

can you teach me how to solve that?

OpenStudy (anonymous):

@Dexter810

OpenStudy (anonymous):

yes :) See , the chain rule is applied when you are differentiating with something other than x ..like x^2 or say , 5x+3 .

OpenStudy (anonymous):

Now , you differentiate your function with respect to the other function that is not x , multiply it with the derivative of ' the other function that is not x' with repect to x

OpenStudy (anonymous):

I hope you undertood what i am trying to say , give it one shot :)

OpenStudy (anonymous):

[ 1/2 sin4x ] ' = 1/2 cos4x * 4 ?

OpenStudy (anonymous):

The product rule is unnecessary here. Use the sine double angle formula: sin (2u) = 2 sin u cos u. Thus y(x) = (1/2) sin(4x). y'(x) = (1/2) cos(4x) * 4 = 2 cos (4x).

OpenStudy (anonymous):

^ @zephy thats exactly what was done and suggested .

OpenStudy (anonymous):

@zephy @Dexter810 can you teach me . how it become y = (1/2) sin 4x

OpenStudy (anonymous):

Yes @jayryan021 ! Exactly :)

OpenStudy (anonymous):

You can derieve that using the first principle , the proof can be found in any book :)

OpenStudy (anonymous):

okay thanks for answering my question :)

OpenStudy (anonymous):

You can watch it from this video too : https://www.youtube.com/watch?v=ojNz3Yd-mnI No problem :)